OpenStudy (fools101):

Tony purchased a rare 1959 Gibson Les Paul guitar in 2000 for $14 thousand. Experts estimate that its value will increase by 12% per year. If this model remains true, when will the value of the guitar be worth $60 thousand? A) 2002 B) 2013 C) 2008 D) 2012 Tony purchased a rare 1959 Gibson Les Paul guitar in 2000 for $14 thousand. Experts estimate that its value will increase by 12% per year. If this model remains true, when will the value of the guitar be worth $60 thousand? A) 2002 B) 2013 C) 2008 D) 2012 @Mathematics

OpenStudy (fools101):

please help!!

OpenStudy (anonymous):

B

OpenStudy (anonymous):

\[60000=(14000)(1.12)^t\]\[4.28=1.12^t\]\[\log_{10} 4.28/\log_{10}1.12=t\]\[t=12.8\]So, yeah, B

OpenStudy (fools101):

Thank u!!!! SO MUCHHHH!!

OpenStudy (anonymous):

np :)

OpenStudy (fools101):

^ ^.!

OpenStudy (anonymous):

Actually given our debate on the other question...the answer here shopuld be 2012. In t=12.8 years the value is $60,000 which is within the year 2012. Pick D